# AGITB: A Signal-Level Benchmark for Evaluating Artificial General Intelligence

### v6

@@ 5 Conclusion
Unlike conventional benchmarks that focus on high-level task performance—such as question answering or language translation—AGITB assesses whether a system exhibits behaviors thought to reflect core operational principles of the biological cortex. It focuses on low-level computational properties that are biologically grounded and essential for the emergence of general intelligence. The proposed testbed introduces a systematic framework comprising twelve fundamental tests that evaluate a model’s ability to learn adaptively from raw input.
AGITB requires models to begin from an unbiased initial state and to develop functionality solely through exposure to structured or unstructured input. This design reflects key insights from contemporary neuroscience regarding input-driven learning and cortical plasticity. In biological systems, high-level reasoning emerges not from symbolic manipulation but from the adaptive prediction of low-level sensory signals. This predictive process is more than pattern matching—it involves the gradual construction of signal-grounded knowledge that enables abstraction and generalization over time.
@@ Appendix A LLM prompt
Your task is to simulate a model that **predicts the next 10-bit input**. The model has an internal state that only updates after receiving an input. Predictions must always be exactly 10 bits long, and the internal state can retain information from hundreds of past inputs. Rules:
@@
    • Update the model’s internal state.
    • Generate a 10-bit prediction based on the model’s state.
@@
    • The model should continuously learn and adapt.

### v7

@@ 5 Conclusion
Unlike conventional benchmarks that focus on high-level task performance—such as question answering or language translation—AGITB assesses whether a system exhibits behaviours thought to reflect core operational principles of the biological cortex. It focuses on low-level computational properties that are biologically grounded and essential for the emergence of general intelligence. The proposed testbed introduces a systematic framework comprising twelve fundamental tests that evaluate a model’s ability to learn adaptively from raw input.
AGITB requires models to begin from an unbiased initial state and to develop functionality solely through exposure to structured or unstructured input. This design reflects key insights from contemporary neuroscience regarding input-driven learning and cortical plasticity. In biological systems, high-level reasoning emerges not from symbolic manipulation but from the adaptive prediction of low-level sensory signals. This predictive process is more than pattern matching—it involves the gradual construction of signal-grounded knowledge that enables abstraction and generalisation over time.
@@
// TODO While individual AGITB requirements may seem trivial in isolation, the central challenge lies in designing a universal AGI architecture that can satisfy all required tasks across the full suite within a cohesive and unified framework. As a recently proposed benchmark, AGITB has not yet been passed by any non-AGI system, and it remains an open question whether such a solution is even theoretically possible.
@@ Appendix A LLM prompt
Your task is to simulate a model that **predicts the next 10-bit input**. The model has an internal state that only updates after receiving an input. Predictions must always be exactly 10 bits long, and the internal state can retain information from hundreds of past inputs.
Rules:
@@
    • Update the internal state of the model.
    • Generate a 10-bit prediction based on the state of the model.
@@
    • The model should learn and adapt continuously.